Eunice is thrice as old as Jason. Four years from now, Eunice will only be twice as old as Jason. How old is Eunice ?
Answer by greenestamps(13200)



x = Jason now
3x = Eunice now

x+4 = Jason 4 years from now
3x+4 = Eunice 4 years from now

4 years from now, Eunice will be twice as old as Jason:

3x+4 = 2(x+4)

Solve using basic algebra....